OpenStudy (anonymous):

A quiz has 1 true-false question and 2 multiple choice questions that each have 4 answers. What is the probability that someone guessing all 3 answers will get 100% on the quiz?

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{4}\times \frac{1}{4}=\frac{1}{32}\]\]